OpenStudy (anonymous):

Find the values of x in this equation: x-15/x=2

OpenStudy (anonymous):

by simplyfying it you get \[x^{2}-2x-15=0 =>x^{2}-5x+3x-15=0 => x(x-5)+3(x-5)=0 =>(x-5)(x+3)=0 =>x=5,-3

OpenStudy (anonymous):

thank youuu

OpenStudy (anonymous):

you welcome//